# AI-Based Vulnerability Analysis of NFT Smart Contracts

### v1

@@ Abstract.
In the research experiment of this article, our research work is divided into several stages. Firstly, we collected a large number of smart contract codes and classified them, identifying several common defects, including Risky Mutably Porxy, ERC-721 Recentrancy, Unlimited Mining, Missing Requirements, and Public Burns. Secondly, we used Python to process the smart contracts. On the one hand, we modified the file names, and on the other hand, we batched the process of the content for analysis and application. Next, we built a model of the decision tree. Firstly, we carried out the feature extraction. We selected the algorithm and divided the data. After comparing and processing, we chose the CART classification tree to process. By gene coefficient, we analyzed and sorted the data, and got the initial model of the decision tree. Then, we introduced the random forest model on the basis of the decision tree. From abstracting the same amount of samples to selecting features randomly.From adjusting and optimizing parameters to completing the construction of the forest model. Finally, we compared and analyzed the decision tree, random forest, and self-built model in the paper and drew general conclusions.
@@ 1. INTRODUCTION
This paper focuses on the currently widely used NFT smart contracts, conducting a comprehensive and in-depth study on their related security issues, collecting a large number of applied NFT smart contract codes; after careful sorting and differentiation, the attacks are divided into five categories; they are processed and studied separately and finally trained, based on a comprehensive accuracy rate of 89.4% of the random forest-based smart contract detection model; in addition, this paper also continues to explore its shortcomings and possible future work directions and priorities, providing its own opinions for the improvement of NFT smart contract detection technology.

### v2

@@ Abstract.
With the rapid growth of the NFT market, the security of smart contracts has become crucial. However, existing AI-based detection models for NFT contract vulnerabilities remain limited due to their complexity, while traditional manual methods are time-consuming and costly. This study proposes an AI-driven approach to detect vulnerabilities in NFT smart contracts.
We collected 16,527 public smart contract codes, classifying them into five vulnerability categories: Risky Mutable Proxy, ERC-721 Reentrancy, Unlimited Minting, Missing Requirements, and Public Burn. Python-processed data was structured into training/test sets. Using the CART algorithm with Gini coefficient evaluation, we built initial decision trees for feature extraction. A random forest model was implemented to improve robustness through random data/feature sampling and multitree integration. GridSearch hyperparameter tuning further optimized the model, with 3D visualizations demonstrating parameter impacts on vulnerability detection.
Results show the random forest model excels in detecting all five vulnerabilities. For example, it identifies Risky Mutable Proxy by analyzing authorization mechanisms and state modifications, while ERC-721 Reentrancy detection relies on external call locations and lock mechanisms. The ensemble approach effectively reduces single-tree overfitting, with stable performance improvements after parameter tuning. This method provides an efficient technical solution for automated NFT contract detection and lays groundwork for scaling AI applications.
@@ 1. INTRODUCTION
This paper focuses on the currently widely used NFT smart contracts, conducting a comprehensive and in-depth study on their related security issues, collecting a large number of applied NFT smart contract codes; after careful sorting and differentiation, the attacks are divided into five categories; they are processed and studied separately and finally trained, in addition, this paper also continues to explore its shortcomings and possible future work directions and priorities, providing its own opinions for the improvement of NFT smart contract detection technology.